A part of $7000 was invested at 4% annual interest and remaining at 9% annual interest. At the end of the year the total amount recieved was $7292. How much money was invested at each rate?
Answer by richwmiller(17219)

In the future do not submit similar problems.
Similar means they are the same type problem but differ by the numbers or addition vs subtraction.
x+y=7000,
0.04*x+0.09*y=292
x=7000-y
0.04*(7000-y)+0.09*y=292
280-0.04y+0.09*y=292
0.05*y=12
y=240 at 9%
x=7000-y
x=6760 at 4%
check
0.04*6760+0.09*240=292
270.4+21.6=292
292=292
ok